OpenStudy (anonymous):

If the simple interest on $6000 for 5 years is $2100, then what is the rate?

OpenStudy (anonymous):

y=yo e^(kt) y=6000+2100 yo=6000 t: time k rate

OpenStudy (anonymous):

Did you figure it out? I have a slightly more simple method

OpenStudy (anonymous):

No, please help.

OpenStudy (anonymous):

Okay. Well I would divide both by 5, that way you can see what interest is earned annually. So 6000/5=1200 and 2100/5=420. So now we know that 420 is 35% of 6000.